# Cost-Efficient Online Decision Making: A Combinatorial Multi-Armed Bandit Approach

### v2

@@ 5 Theoretical Analysis
In this section, we theoretically analyze the framework proposed in Section 4. Specifically, we provide an upper bound on the Bayesian regret (defined below) of the online decision making procedure in Algorithm 1.
In our problem setting, the set of feasible super arms $\mathcal{I}^{(t)}$ is restricted such that $y^{(t)}=y^{*}$ (i.e., we study the setting where correct decisions are guaranteed, and the objective is to minimize the cost of tests), and we assume that we have access to a perfect oracle which can solve the optimization problem Eq. 4.
The Bayesian regret is defined as the expected value of $\text{Regret}(T)$ (see Eq. 3), with the expectation taken over the prior distribution of $\boldsymbol{\mu}^{*}$ and other random variables (e.g., randomness in the rewards and the policy of the agent).
@@
We begin our regret analysis by presenting a decomposition of $\mathbb{E}[\Delta^{(t)}_{\boldsymbol{\mu}^{*}}(S^{(t)})]$ in the following lemma.
@@
For any upper confidence bound $U^{(t)}:\mathcal{I}^{(t)}\rightarrow\mathbb{R}$ and lower confidence bound $L^{(t)}:\mathcal{I}^{(t)}\rightarrow\mathbb{R}$ we have
@@
We give the full proof, which is based on Proposition 1 of Russo and Van Roy (2014), in Appendix A. ∎
Additionally, we define lower and upper confidence bounds as follows:
@@
where $\hat{\mu}^{(t)}_{ij}$ is the average cost of base arm $a_{ij}$ till time $t$, and $N^{(t)}(ij)$ is the number of times base arm $a_{ij}$ has been played till time $t$. We continue by bounding each term in Eq. 13.
@@
To prove Lemma 5.3, we use Eq. 1 which implies that $|\mu^{*}_{ij}-\hat{\mu}^{(t-1)}_{ij}\big{|}=|\mu_{ij}^{(1)}-\mu_{ij}^{(0)}||\theta_{ij}^{*}-\hat{\theta}_{ij}^{(t-1)}|$ together with the fact that $|\mu_{ij}^{(1)}-\mu_{ij}^{(0)}|\leq 1$, as well as Hoeffding’s inequality. The full proof is given in Appendix A. ∎
Continuing from Eq. 16, using Lemma 5.3, and noting that $\big{|}\mu^{*}_{ij}-\hat{\mu}^{(t-1)}_{ij}\big{|}\leq 1$ we have
@@
In order to bound the last term of the regret decomposition, we utilize the following lemma.
Lemma 5.4.
Proof.
We prove Lemma 5.4 using a similar technique as Lemma 8 by Åkerblom et al. (2023) for the combinatorial batched feedback setting (with batch size 1). See Appendix A for the full proof. ∎
Now, based on Eq. 17, Eq. 18, Eq. 13, and Lemma 5.4 we have
Bayesian $\text{Bayesian Regret}(T)\leq 4mn+2mn\sqrt{8T\log T}=\mathcal{O}(mn\sqrt{T\log T}).$ (19)
We utilize the structure of the problem in our analysis and thereby, the bound derived in Theorem 5.1 is significantly tighter than the ones by Rahbar et al. (2023); Chen et al. (2017) which take a different perspective than our CMAB formulation. Our bound is also consistent with Bayesian upper bounds for combinatorial semi-bandit methods developed in other settings.
[…]
6 Experiments
In this section, we empirically validate our cost-efficient online decision making framework via various experiments. Unless otherwise specified, we consider Beta(2,2) as the prior distribution of all $\theta_{ij}^{*}$’s. We employ both of the cost-efficient approximate oracles introduced in Section 4.1 (W-IG and W-EC2) as the oracle in Algorithm 1. We also employ the hypothesis enumeration procedure proposed in Chen et al. (2017) to generate the most likely hypotheses for each decision region. In Appendix B, we demonstrate the applicability of our framework in the setting where the decision regions for the set of hypotheses are not known. Moreover, in Section 6.2, we provide an extension of our framework to real-valued non-binary test outcomes. We run each experiment with 5 different random seeds.
Data sets.
We apply our framework to the LED display domain data set from the UCI repository (Breiman et al., 1988). Additionally, we use the ProPublica recidivism (Compas) data set (Larson et al., 2016) and the Fair Isaac (Fico) credit risk data set (FICO et al., 2018), which both are preprocessed in the same way as by Hu et al. (2019). The aforementioned data sets contain relatively large numbers of tests and few decision outcomes, which correspond well to, e.g., the medical diagnosis example in Section 1. In applications like the charging station selection example, it might instead be reasonable to expect a large number of decision outcomes (individual charging stations or groups sharing certain characteristics) and relatively few tests (questions). Hence, in order to thoroughly evaluate the latter type of setting, we create a synthetic data set (called Navigation) with the desired characteristics. With $n=5$ tests and $m=20$ decision outcomes, we sample parameters $\theta_{ij}^{*}$ from the prior distribution Beta(2,2) for each $i\in[n]$ and $j\in[m]$, which subsequently use to generate the data set. Moreover, to demonstrate the applicability of our methods to real-valued test outcomes, we investigate the Breast Cancer Wisconsin data set (Street et al., 1993) (Section 6.2). Finally, in Section 6.3, we present the application of our framework to a real-world troubleshooting case study.
Algorithms.
  • DPP. Determinantal Point Process (DPP) provides a way to select diverse sets. With this method, we use exact sampling (Derezinski et al., 2019) (finite DPP with likelihood kernel) to subsample columns (tests). We use the implementation of Gautier et al. (2019). To perform the sampling, in each time step $t$, we provide all data points (all test results for all data points) until time $t-1$ to the sampler. The number of tests selected by the sampler is related to the number of eigenvectors of the kernel matrix.
6.1 Experimental Results
In Figure 1, we illustrate the cost of performing tests (i.e., $-\mathbb{E}[R(S^{(t)})]$) during online learning for different algorithms. The cost of each test is calculated based on Eq. 1, where the parameter $\theta_{ij}$ is derived from the data. Our results in Figure 1 show that our framework yields the lowest information acquisition cost for all data sets. DPP has a low cost initially since it only uses a few data points to perform sampling, but after a few time steps its cost becomes as high as ‘All’. Additionally, we observe that the W-EC2 algorithm always has the lowest cost. This result is consistent with the theoretical results of Golovin et al. (2010) for the cost of the original EC2 algorithm in the offline setting. We also observe that the exploration methods (TS and BayesUCB) exhibit very similar performances in terms of cost.
During online learning, it is observed that the cost of the random algorithm remains relatively constant, whereas the costs associated with the W-IG and W-EC2 algorithms decrease until they converge to a low value. This behavior is attributed to the utilization of parameter estimates by the W-IG and W-EC2 algorithms when selecting tests. As the knowledge of the agent about these parameters improves over the course of learning, the associated costs decline.
Furthermore, in Table 1, we report the average cost of decision making in a single time step for different data sets. It is obvious that our framework incurs significantly lower costs for performing tests.
6.2 Extension to Real-Valued Test Outcomes
In this section, we extend our experimental results to real-valued (non-binary) test outcomes. For this purpose, we adapt the discretization method proposed by Rahbar et al. (2023). Specifically, for each test, we consider different thresholds for “binarization”. Then, in each time step, we can calculate the gain ($\Delta_{\text{W-IG}}$ or $\Delta_{\text{W-EC}^{2}}$) for the tests using different binarization thresholds and choose the one with a maximal gain. We employ this method on the Breast Cancer Wisconsin data set. In this data set, we predict the diagnosis from 30 real-valued medical tests computed from an image of a fine needle aspirate. Figure 2 illustrates the decision making cost incurred by our framework when applied to this data set. Similar to the results depicted in Figure 1, our framework consistently achieves the lowest cost, demonstrating its effectiveness in this setting. In particular, we observe that W-EC2-TS yields the best results, and W-IG-TS, W-EC2-BUCB, and W-IG-BUCB are other suitable methods for this task.
6.3 Application to Online Troubleshooting
We experiment on approximately 1500 different scenarios (i.e., roughly 100 scenarios for each decision). A scenario starts with a customer entering the system. Then the online learning agent aims at finding the correct decision by asking a sequence of multiple questions of the customer (performing tests). The goal for the agent is to ask the most informative questions in a cost-efficient manner.
Figure 3 shows the cost of making decisions for different time steps (i.e., scenarios) in the same setting as Section 6.1. The results show that our framework enables the agent to find the correct decisions for troubleshooting in mobile devices with a significantly lower cost compared to other methods. Similar to previous observations, the W-EC2 algorithm remains the most cost-effective method within our framework.
7 Conclusion
[…]
Acknowledgments
Appendix A Additional Proofs
A.1 Proof of Lemma 5.2
Proof.
This lemma is based on Proposition 1 of Russo and Van Roy [2014]. Given the history of played super arms (and base arms) and their rewards, we know that in Thompson Sampling the optimal super arm $S^{*,(t)}$ and the played super arm $S^{(t)}$ follow the same distribution. Since $U^{(t)}(\cdot)$ is a deterministic function, we conclude that given the history, the expected values of $U^{(t)}(S^{(t)})$ and $U^{(t)}(S^{*,(t)})$ are equal, and we have:
$\mathbb{E}[\Delta^{(t)}_{\boldsymbol{\mu}^{*}}(S^{(t)})]=\mathbb{E}[U^{(t)}(S^{(t)})-f_{\boldsymbol{\mu}^{*}}(S^{(t)})]+\mathbb{E}[f_{\boldsymbol{\mu}^{*}}(S^{*,(t)})-U^{(t)}(S^{*,(t)})],$
and by adding and subtracting $\mathbb{E}[L^{(t)}(S^{(t)})]$ we prove 13. ∎
A.2 Proof of Lemma 5.3
@@ Appendix A Additional Proofs
=$\begin{split}\stackrel{{\scriptstyle Eq.6}}{{=}}Pr\Bigg{\{}\exists t\in[T]\;\exists ij\in\mathcal{A},\big{|}\mu_{ij}^{(1)}-\mu_{ij}^{(0)}\big{|}\big{|}\theta_{ij}^{*}-\hat{\theta}_{ij}^{(t-1)}\big{|}>\sqrt{\frac{2\log T}{N^{(t-1)}(ij)}}\Bigg{\}}\end{split}$
@@
The last inequality is due to Hoeffding’s inequality (e.g., Theorem A.1 of Slivkins et al. [2019] with $\alpha=2$).
A.3 Proof of Lemma 5.4
@@
Appendix B Additional Experimental Results
In this section, to demonstrate the generality of our framework, we examine the setting where the decision regions for the enumerated hypotheses are not known. This situation may lead to incorrect decisions by the agent. Specifically, in the context of online learning, our objectives are: i) to achieve highly accurate decisions, and ii) to minimize the cost of performing tests. To quantify the accuracy of the decisions, we employ a utility function. Specifically, we assign a utility of $2$ for a correct decision and $-1$ for an incorrect one. Additionally, we consider an “unknown" decision when the correct hypothesis is assigned to multiple decision regions. We define the utility of such decisions as 0.
The upper row in Figure 4 shows the cost of making decisions for different data sets (when the correct decisions of the hypotheses are not known). Similar to the findings in Figure 1, we observe that our framework (with both W-IG and W-EC2) yields a final decision with significantly lower cost than other algorithms. Again, we observe that EC2 (with both Thompson Sampling and BayesUCB) has the lowest cost of performing tests. The IG algorithm results in the second lowest cost, and the random algorithm for picking tests is the third option.
As previously mentioned, when faced with unknown decision regions, our goal is to enhance the accuracy of our decisions during online learning. The lower row in Figure 4 illustrates the utility achieved by various information acquisition algorithms. Notably, our framework demonstrates comparable performance to both All and DPP in the early stages of learning. Therefore, our approach proves capable of making precise decisions with a significantly reduced cost. We note that DPP yields very low costs in the early time steps, but the corresponding utility is also very low.
Figure 4: Decision making cost (upper row) and utility (lower row) of our framework applied to the setting that the decision regions are not known beforehand.

### v3

@@ 5 Theoretical Analysis
In this section, we theoretically analyze the framework proposed in Section 4. Specifically, we provide an upper bound on the Bayesian regret (defined below) of the online decision making procedure in Algorithm 2. The analysis technique that we use extends the approach by Russo & Van Roy (2014) to our setting and yields an analysis which is significantly simpler than those by Chen et al. (2013) and Wang & Chen (2018). Furthermore, in our problem, due to the interactive nature of the oracle, the set of available arms changes in each time step. In general, a bandit problem with changing sets of available arms is called a bandit with sleeping arms, a setting which is not considered in the analyses of combinatorial UCB (CUCB) by Chen et al. (2013) and combinatorial Thompson sampling (CTS) by Wang & Chen (2018). Our analysis holds for interactive oracles through the sleeping arms assumption. Moreover, the analyses of CUCB and CTS yield instance-dependent regret bounds, whereas we focus on an instance-independent (and Bayesian) bound, which also requires a different type of analysis (e.g., a different regret decomposition). Furthermore, the analysis of CTS is tailored for one specific choice of uninformative prior, while ours, in principle, allows for an arbitrary selection of the prior distribution (whether uninformative or informative).
In our problem setting, the set of feasible super arms $\mathcal{I}^{(t)}$ is restricted such that $y^{(t)}=y^{*}$ (i.e., correct decisions are guaranteed, and the objective is to minimize the cost of tests). This means that a super arm is feasible if and only if the test outcomes determine the correct decision region.  [^4] We relax this assumption in the experiments of Section 6.4. We also consider access to a perfect oracle which can solve the optimization problem Eq. 4. As mentioned earlier, the problem under consideration is NP-hard. On the other hand, assuming access to an exact oracle is common for analyses of regret bounds for CMABs, even in settings where the exact oracle needs to deal with an NP-hard problem. For instance, the work of Hüyük & Tekin (2020) assumes access to exact oracles for NP-hard network problems. Åkerblom et al. (2023) consider the intractable problem of stochastic bottleneck identification in an online setting. In fact, via the elegant CMAB formulation of our problem (in contrast to, e.g., the POMDP formulation of Rahbar et al. 2023 and Chen et al. 2017), we are able to obtain a regret bound that is linear in the number of tests (rather than exponential).
The Bayesian regret is defined as the expected value of $\text{Regret}(T)$ (see Eq. 3), with the expectation taken over the prior distribution of $\boldsymbol{\mu}^{*}$ and other random variables (e.g., randomness in the rewards and the policy of the agent). Formally,
@@
We begin our regret analysis by a decomposition of $\mathbb{E}[\Delta^{(t)}_{\boldsymbol{\mu}^{*}}(S^{(t)})]$ in the following lemma.
@@
For any upper confidence bound $U^{(t)}:\mathcal{I}^{(t)}\rightarrow\mathbb{R}$ and lower confidence bound $L^{(t)}:\mathcal{I}^{(t)}\rightarrow\mathbb{R}$,
@@
This lemma is based on Proposition 1 of Russo & Van Roy (2014). Given the history of played super arms (and base arms) and their rewards, we know that in Thompson Sampling the optimal super arm $S^{*,(t)}$ and the played super arm $S^{(t)}$ follow the same distribution. Since $U^{(t)}(\cdot)$ is a deterministic function, we conclude that given the history, the expected values of $U^{(t)}(S^{(t)})$ and $U^{(t)}(S^{*,(t)})$ are equal, and we have:
$\begin{split}\mathbb{E}[\Delta^{(t)}_{\boldsymbol{\mu}^{*}}(S^{(t)})]=\mathbb{E}[U^{(t)}(S^{(t)})-f_{\boldsymbol{\mu}^{*}}(S^{(t)})]+\mathbb{E}[f_{\boldsymbol{\mu}^{*}}(S^{*,(t)})-U^{(t)}(S^{*,(t)})],\end{split}$
and by adding and subtracting $\mathbb{E}[L^{(t)}(S^{(t)})]$ we prove Lemma 5.2. ∎
Additionally, we define lower and upper confidence bounds as
@@
where $\hat{\mu}^{(t)}_{ij}$ is the average cost of base arm $a_{ij}$ till time $t$, and $N^{(t)}(ij)$ is the number of times base arm $a_{ij}$ has been played till time $t$. We continue the proof by bounding each term in Lemma 5.2. For the second and third terms, we have the following lemma (where the proof is in Appendix B.1).
Lemma 5.3.
$\mathbb{E}[L^{(t)}(S^{(t)})-f_{\boldsymbol{\mu}^{*}}(S^{(t)})]\leq\frac{2mn}{T},\mathbb{E}[f_{\boldsymbol{\mu}^{*}}(S^{*,(t)})-U^{(t)}(S^{*,(t)})]\leq\frac{2mn}{T}$.
In order to bound the first term of the regret decomposition, we utilize the following lemma (where the proof uses a similar technique as Lemma 8 by Åkerblom et al. (2023) for the combinatorial batched feedback setting (with batch size 1). See Appendix B.3).
Lemma 5.4.
Now, based on Lemma 5.2, Lemma 5.3, and Lemma 5.4 we have Bayesian $\text{Bayesian Regret}(T)\leq 4mn+2mn\sqrt{8T\log T}=\mathcal{O}(mn\sqrt{T\log T})$. ∎
We utilize the structure of the problem in our analysis and thereby, the bound derived in Theorem 5.1 is significantly tighter than the ones by Rahbar et al. (2023); Chen et al. (2017) which take a different perspective than our CMAB formulation. Our bound is also consistent with Bayesian upper bounds for combinatorial semi-bandit methods developed in other (standard) settings.
6 Experiments
In this section, we empirically validate our cost-efficient online decision making framework via various experiments  [^5] The source code for our experiments can be accessed here on GitHub. . Unless otherwise specified, we consider Beta(2,2) as the prior distribution of all $\theta_{ij}^{*}$’s. We employ both of the cost-efficient approximate oracles introduced in Section 4.1 (W-IG and W-EC2) as the oracle in Algorithm 2. We also employ the hypothesis enumeration procedure proposed by Chen et al. (2017) to generate the most likely hypotheses for each decision region. In Section 6.3, we provide an extension of our framework to real-valued non-binary test outcomes. Moreover, In Section 6.4, we demonstrate the applicability of our framework in the setting where the decision regions for the set of hypotheses are not known. We run each experiment with 5 different random seeds on a single-node CPU machine with 8 cores, 16GB memory and macOS. We implement the algorithms in Python mainly using NumPy (Harris et al., 2020), NetworkX (Hagberg et al., 2008) and SciPy (Virtanen et al., 2020).
Data sets.
We apply our framework to the LED display domain data set from the UCI repository (Breiman et al., 1988). Additionally, we use the ProPublica recidivism (Compas) data set (Larson et al., 2016) and the Fair Isaac (Fico) credit risk data set (FICO et al., 2018), which both are preprocessed in the same way as by Hu et al. (2019). The aforementioned data sets contain relatively large numbers of tests and few decision outcomes, which correspond well to, e.g., the medical diagnosis example in Section 1. In applications like the charging station selection example, it might instead be reasonable to expect a large number of decision outcomes (individual charging stations or groups sharing certain characteristics) and relatively few tests (questions). Hence, in order to thoroughly evaluate the latter type of setting, we create a synthetic data set (called Navigation) with the desired characteristics. With $n=5$ tests and $m=20$ decision outcomes, we sample parameters $\theta_{ij}^{*}$ from the prior distribution Beta(2,2) for each $i\in[n]$ and $j\in[m]$, which subsequently use to generate the data set. Moreover, in Section 6.2, we present the application of our framework to a real-world troubleshooting case study. Finally, to demonstrate the applicability of our methods to real-valued test outcomes, we investigate the Breast Cancer Wisconsin data set (Street et al., 1993) (Section 6.3).
Algorithms.
DPP. Determinantal Point Process (DPP) provides a way to select diverse sets. With this method, we use exact sampling (Derezinski et al., 2019) (finite DPP with likelihood kernel) to subsample columns (tests). We use the implementation of Gautier et al. (2019). To perform the sampling, in each time step $t$, we provide all data points (all test results for all data points) until time $t-1$ to the sampler. The number of selected tests is related to the number of eigenvectors of the kernel matrix.
[…]
6.1 Experimental Results
In Figure 2, we illustrate the cost [^6] The costs reported here directly correspond to the (instant) regret up to only a constant term. Therefore, the order of the methods and their relative performances remain the same. of performing tests (i.e., $-\mathbb{E}[R(S^{(t)})]$) during online learning for different algorithms. The cost of each test is calculated based on Eq. 1, where the parameter $\theta_{ij}$ is derived from the data. The (maximum) numbers of enumerated hypotheses per decision region are 2, 5, 70 and 70 for Navigation, LED, Fico and Compas datasets, respectively. Our results in Figure 2 show that our framework yields the lowest information acquisition cost for all data sets. DPP has a low cost initially since it only uses a few data points to perform sampling, but after a few time steps its cost becomes as high as ‘All’. During these initial time steps DPP can make incorrect decisions. Additionally, we observe that the W-EC2 algorithm always has the lowest cost. This result is consistent with the theoretical results of Golovin et al. (2010) for the cost of the original EC2 algorithm in the offline setting. We also observe that the exploration methods (TS and BayesUCB) exhibit very similar performances in terms of cost. During online learning, it is observed that the cost of the random algorithm remains relatively constant, whereas the costs associated with the W-IG and W-EC2 algorithms decrease until they converge to a low value. This behavior is attributed to the utilization of parameter estimates by the W-IG and W-EC2 algorithms when selecting tests. As the knowledge of the agent about these parameters improves over the course of learning, the associated costs decline. Furthermore, in Table 1, we report the average cost of decision making in a single time step for different data sets. It is obvious that our framework incurs significantly lower costs for performing tests.
6.2 Application to Online Troubleshooting
We experiment on approximately 1500 different scenarios (i.e., roughly 100 scenarios for each decision). A scenario starts with a customer entering the system. Then the online learning agent aims at finding the correct decision by asking a sequence of multiple questions of the customer (performing tests). The goal for the agent is to ask the most informative questions in a cost-efficient manner. We enumerate 15 hypotheses for each decision. Figure 3 shows the cost of making decisions for different time steps (i.e., scenarios) in the same setting as Section 6.1. The results show that our framework enables the agent to find the correct decisions for troubleshooting in mobile devices with a significantly lower cost compared to other methods. Similar to previous observations, the W-EC2 algorithm remains the most cost-effective method within our framework.
Figure 5: Decision making cost (first row) and utility (second row) of our framework applied to the setting that the decision regions are not known beforehand.
6.3 Extension to Real-Valued Test Outcomes
In this section, we extend our experimental results to real-valued (non-binary) test outcomes. For this purpose, we adapt the discretization method proposed by Rahbar et al. (2023). Specifically, for each test, we consider different thresholds for “binarization”. Then, in each time step, we can calculate the gain ($\Delta_{\text{W-IG}}$ or $\Delta_{\text{W-EC}^{2}}$) for the tests using different binarization thresholds and choose the one with a maximal gain. We employ this method on the Breast Cancer Wisconsin data set. In this data set, we predict the diagnosis from 30 real-valued medical tests computed from an image of a fine needle aspirate. Figure 4 illustrates the decision making cost incurred by our framework when applied to this data set. We enumerate 70 hypotheses per decision region. Similar to the results depicted in Figure 2, our framework consistently achieves the lowest cost, demonstrating its effectiveness in this setting. In particular, we observe that W-EC2-TS yields the best results, and W-IG-TS, W-EC2-BUCB, and W-IG-BUCB are other suitable methods for this task.
6.4 Extension to Unknown Decision Regions
In this section, to demonstrate the generality of our framework, we examine the setting where the decision regions for the enumerated hypotheses are not known. To assign hypotheses to decision regions, we employ the hypothesis enumeration algorithm discussed in Section 6.1 to generate the most probable hypotheses for each decision region based on the current parameter sample. Therefore, each hypothesis is assumed to correspond to the decision it is assigned to. This situation may lead to incorrect decisions by the agent. Specifically, in the context of online learning, our objectives are: i) to achieve highly accurate decisions, and ii) to minimize the cost of performing tests. To quantify the accuracy of the decisions, we employ a utility function. Specifically, we assign a utility of $2$ for a correct decision and $-1$ for an incorrect one. Additionally, we consider an “unknown” decision when the correct hypothesis is assigned to multiple decision regions. We define the utility of such decisions as 0.
The first row in Figure 5 shows the cost of making decisions for different data sets (when the correct decisions of the hypotheses are not known). We use the same numbers of enumerated hypotheses as Figure 2. Similar to the findings in Figure 2, we observe that our framework (with both W-IG and W-EC2) yields a final decision with significantly lower cost than other algorithms. Again, we observe that EC2 (with both Thompson Sampling and BayesUCB) has the lowest cost of performing tests. The IG algorithm results in the second lowest cost, and the random algorithm for picking tests is the third option.
As previously mentioned, when faced with unknown decision regions, our goal is to enhance the accuracy of our decisions during online learning. The second row in Figure 5 illustrates the utility achieved by various information acquisition algorithms. Notably, our framework demonstrates comparable performance to both All and DPP in the early stages of learning. Therefore, our approach proves capable of making precise decisions with a significantly reduced cost. We note that DPP yields very low costs in the early time steps, but the corresponding utility is also very low.
7 Conclusion
We propose a novel framework for online decision making based on an elegant design of a combinatorial multi-armed bandit problem, which incorporates the cost of performing tests, and where the costs may be stochastic and depend on both test and decision outcomes. Within this framework, we develop various cost-efficient online decision making methods such as W-EC2 and W-IG. We also adapt Thompson Sampling and BayesUCB, methods that are commonly used for exploration. In particular, we provide a theoretical upper bound for the Bayesian regret of Thompson Sampling. We demonstrate the performance of the framework on a number of data sets from different domains.
Acknowledgments
Appendix A Nomenclature
Table 2: Table of notation used in this paper.
Notation Description
$t$ Time step
$\mathbf{x}$ A problem instance
$\mathbf{x}^{(t)}$ A problem instance received at time $t$
$i$ Test index
$x_{i}$ Test with index $i$
$x_{i}^{(t)}$ Test with index $i$ performed at time $t$
$y$ A decision
$y^{(t)}$ A decision made at time $t$
$n$ Number of available tests
$m$ Number of possible decisions
$y^{*}$ Correct decision
$T$ Time horizon
$a_{ij}$ Base arm related to test with index $i$ and decision with index $j$
$\mathcal{A}$ Set of base arms
$S$ A super arm
$S^{(t)}$ A super arm selected at time $t$
$R^{(t)}$ Reward received at time $t$
$R(S)$ Reward of super arm $S$
$\mathcal{I}^{(t)}$ Set of feasible super arms at time $t$
$\theta_{ij}$ Parameter of Bernoulli distribution for test $i$ and decision $j$
$\mu_{ij}^{(0)}$ Cost of base arm $a_{ij}$ when test outcome is 00
$\mu_{ij}^{(1)}$ Cost of base arm $a_{ij}$ when test outcome is $1$
$\mu_{ij}$ Expected cost of base arm $a_{ij}$
$\boldsymbol{\mu}$ Vector of all $\mu_{ij}$’s
$\boldsymbol{\theta}$ Vector of all $\theta_{ij}$’s
$\boldsymbol{\mu}^{*}$ True mean cost vector
$\boldsymbol{\theta}^{*}$ True parameter vector
$\Delta^{(t)}_{\boldsymbol{\mu}^{*}}(S)$ Suboptimality gap at time $t$ of super arm $S$ with mean cost vector is $\boldsymbol{\mu}^{*}$
$f_{\boldsymbol{\mu}}(S)$ Expected reward of super arm $S$ given a mean cost vector $\boldsymbol{\mu}$
$S^{*,(t)}$ Optimal super arm at time $t$
$\text{Regret}_{\boldsymbol{\mu}^{*}}(T)$ Regret for a time horizon $T$ when true mean cost vector is $\boldsymbol{\mu}^{*}$
$\alpha_{ij}^{(t)}$ First shape parameter of Beta distribution related to $\theta_{ij}$ at time $t$
$\beta_{ij}^{(t)}$ Second shape parameter of Beta distribution related to $\theta_{ij}$ at time $t$
$Q_{\lambda}(\eta)$ Quantile function (if X has distribution $\lambda$, then $\text{Pr}\left(X\leq Q_{\lambda}(\eta)\right)=\eta$)
$\overline{\theta}^{(t)}_{ij}$ Upper confidence bound for $\theta_{ij}$ used in BayesUCB
$\underline{\theta}^{(t)}_{ij}$ Lower confidence bound for $\theta_{ij}$ used in BayesUCB
$\mathcal{H}$ Set of all possible hypotheses
$h$ A hypothesis
$\mathcal{P}$ Set of performed tests
$\mathbf{x}_{\mathcal{P}}$ Vector of results of tests in $\mathcal{P}$
$\mathbb{H}(z)$ Shannon entropy of a random variable $z$
$\Delta_{\text{Alg}}\left(i\mid\mathbf{x}_{\mathcal{P}}\right)$ Gain of test $i$ with algorithm Alg when $\mathcal{P}$ is the set of performed tests
$i^{*}$ Index of next test to perform
$\mathcal{H}_{i}$ Set of hypotheses consistent with the outcome of test $i$
$(h,h^{\prime})$ An edge between two hypotheses $h$ and $h^{\prime}$
$w_{hh^{\prime}}$ Weight of an edge $(h,h^{\prime})$
$\mathcal{E}$ A set of edges
$w(\mathcal{E})$ Weight of a set of edges $\mathcal{E}$
$K(i)$ Set of edges cut by test $i$
$f_{\text{EC}^{2}}\left(\mathcal{P}\right)$ Objective function of EC2 for a set of performed tests $\mathcal{P}$
Bayesian $\text{Bayesian Regret}(T)$ Bayesian regret over a time horizon $T$
$U^{(t)}(S)$ Upper confidence bound for a super arm $S$
$L^{(t)}(S)$ Lower confidence bound for a super arm $S$
$\hat{\mu}^{(t)}_{ij}$ Average cost of base arm $a_{ij}$ till time $t$
$N^{(t)}(ij)$ Number of times base arm $a_{ij}$ has been played till time $t$
$\bar{x}_{ij}^{(\tau)}$ Average of first $\tau$ observations of test $i$ under decision $j$
Appendix B Additional Proofs
B.1 Proof of Lemma 5.3
Proof.
@@ Appendix B Additional Proofs
To prove Lemma B.1, we use Eq. 1 which implies that $|\mu^{*}_{ij}-\hat{\mu}^{(t-1)}_{ij}\big{|}=|\mu_{ij}^{(1)}-\mu_{ij}^{(0)}||\theta_{ij}^{*}-\hat{\theta}_{ij}^{(t-1)}|$ together with the fact that $|\mu_{ij}^{(1)}-\mu_{ij}^{(0)}|\leq 1$, as well as Hoeffding’s inequality. The full proof is given in Appendix B.2. ∎
Continuing from Eq. 16, using Lemma B.1, and noting that $\big{|}\mu^{*}_{ij}-\hat{\mu}^{(t-1)}_{ij}\big{|}\leq 1$ we have
@@
B.2 Proof of Lemma B.1
@@
[…]
@@
The last inequality is due to Hoeffding’s inequality (e.g., Theorem A.1 of Slivkins et al. (2019) with $\alpha=2$).
B.3 Proof of Lemma 5.4